OpenStudy (anonymous):

How much would you need to deposit in an account in order to have $20,000 in the account in 4 years? Assume the account earns 5% interest.

OpenStudy (anonymous):

do 20,000 x 20% (.2) = 4000. now 20,000-4,000 = 16,000 why did i do 20%? because 5%x4 = 20%

OpenStudy (anonymous):

please correct me if I'm wrong, @parthkohli

OpenStudy (anonymous):

Sounds good to me... I'll let you know. Thank you.

OpenStudy (anonymous):

\[p_{N}=p _{o}(1+\frac{ r }{ k })^{Nk}\] this is the formula for it I just cant find the unknown

OpenStudy (anonymous):

How is the interest added in... does it compound annually, quarterly, monthly, daily??

OpenStudy (anonymous):

Thats what I dont get. It doesn't say, so I can only assume that it is yearly. I brought that up to my professor but have not received a responds yet.

OpenStudy (anonymous):

Thought maybe someone on here could help me out :/

OpenStudy (anonymous):

well if it is yearly then you earn 1200 a year with an intial investment of 2000. Just try out several differnt starting points and calculate the interest until you get 20000

OpenStudy (anonymous):

I'll try it again, and give you my answer and see if you think its correct or not... thx

OpenStudy (anonymous):

if you start with 5882 you are a little bit too short, if you start with 5883 you are a little bit to much. The interest can't compound yearly.

OpenStudy (anonymous):

Also is it just one deposit or do you deposit a set amount and then more each month?

OpenStudy (anonymous):

It is just a one time deposit

OpenStudy (anonymous):

You need more information... im taking a class where all we do are problems like this but we get more information.

OpenStudy (anonymous):

If I knew how to use that formula I could solve the problem.

OpenStudy (anonymous):

So would I try different compounds? .... \[(1+\frac{ 0.05 }{ 4 })^{4}\] this is how i would set up the last part but I just cant figure out the first part. Im taking a math 142 its a teachers math class.... but im horrible at math already

OpenStudy (anonymous):

Has your teacher messaged you back yet?

OpenStudy (anonymous):

Its an online class, so it takes a while for him to respond

OpenStudy (anonymous):

I found out that it is supposed to be compounded monthly....

OpenStudy (amistre64):

Fv = Pv(1+r)^(n); where Fv is the future value, Pv the present value, r is the interest per period, and n is the number of periods in which it gets calculated. interest rate per period is simply take the year amount and divide by 12 a month is a period

OpenStudy (anonymous):

Is this using the formula above? or is this something else?

OpenStudy (amistre64):

same basic idea, but a simpler construct. there are 12*4 periods; with an interest rate of .05/12 per period so its the same content just a different way of presenting it.

OpenStudy (amistre64):

i got to head to an exam, ill try to be back later if the library is open

OpenStudy (anonymous):

oh ok thank you, ill try it again. Ive been stuck on this problem. Good luck! thanks

OpenStudy (anonymous):

\[P_{4}=P_{0}\left( 1+\frac{ 0.05 }{ 12 } \right)^{12(4)}=24,417.91-20,000=4,417.91\] Does this look correct? if not could you let me know where im wrong

OpenStudy (amistre64):

\[20,000=P_o(1+\frac{.05}{12})^{12(4)}\] \[20,000(1+\frac{.05}{12})^{-12(4)}=P_o\] that looks better to me since we know the future value and want to determine the present value.